OpenStudy (anonymous):

Solve the equation -6+3x= -9

OpenStudy (hockeychick23):

ok so first we take your equation -6+3x= -9 and add 6 to both sides to get 3x alone: -6+6+3x=-9+6 Now we have: 3x= -3 Now we divide both sides by 3 to isolate the variable: 3x/3=-3/3 x=-1

OpenStudy (anonymous):

Thank you

OpenStudy (hockeychick23):

Anytime :)

OpenStudy (anonymous):

Yup it is right